OpenStudy (anonymous):

divide (8x^4y3+4x^3y^2-2x^2y-12x^2y^4) by -2x^2y justify each step

OpenStudy (noelgreco):

What's to divide?

OpenStudy (ranga):

The problem is incomplete. It is usually divide something by something. You have not specified what to divide by.

OpenStudy (anonymous):

divide (8x^4y3+4x^3y^2-2x^2y-12x^2y^4) by -2x^2y

OpenStudy (anonymous):

wooow what?? sorry im not good at math

OpenStudy (anonymous):

heres the whole problem 5.) Part A: divide (8x^4y3+4x^3y^2-2x^2y-12x^2y^4) by -2x^2y justify each step Part B: how would your answer in part A be affected if the x^2 variable in the denominator was just an x? Part C: what are the degree and classification of the polynomial you got in part A.

OpenStudy (ranga):

8x^4y^3+4x^3y^2-2x^2y-12x^2y^4 = 2x^2y(4x^2y^2 + 2xy - 1 - 6y^3) Divide 2x^2y(4x^2y^2 + 2xy - 1 - 6y^3) by -2x^2y -4x^2y^2 - 2xy + 1 + 6y^3

OpenStudy (ranga):

B) If the denominator is changed from x^2 to x, it would be equivalent to multiplying the numerator by x. And so the answer to A) should be multiplied by x: -4x^3y^2 - 2x^2y + x + 6xy^3

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

@ranga in part a did you simplify the equation before we multiplied?

OpenStudy (mertsj):

\[\frac{8x^4y^3+4x^3y^2-2x^2y-12x^2y^4}{-2x^2y}=\] \[\frac{8x^4y^3}{-2x^2y}+\frac{4x^3y^2}{-2x^2y}-\frac{2x^2y}{-2x^2y}-\frac{12x^2y^4}{-2x^2y}=\] \[-4x^2y^2-2xy+1+6y^3\]

OpenStudy (anonymous):

@Mertsj are you answering part a

OpenStudy (anonymous):

because @ranga already answered that one

OpenStudy (anonymous):

part a: -4x^2y^2 - 2xy + 1 + 6y^3 part b:If the denominator is changed from x^2 to x, it would be equivalent to multiplying the numerator by x. And so the answer to A) should be multiplied by x: -4x^3y^2 - 2x^2y + x + 6xy^3 part C: third degree polynomial

OpenStudy (anonymous):

@Mertsj are those answers right?

OpenStudy (mertsj):

b is ok c is degree 4

OpenStudy (anonymous):

i thought the degree was the highest exponent @Mertsj

OpenStudy (mertsj):

To find the degree of a polynomial or monomial with more than one variable for the same term, just add the exponents for each variable to get the degree Degree of x^3y^2. Degree of this monomial = 3 + 2 = 5

OpenStudy (anonymous):

are you sure?

OpenStudy (anonymous):

@Mertsj i really feel like its a third degree

OpenStudy (mertsj):

This is America. You have a right to be wrong. And I might point out that the correct answer to a math problem is not based on how you "feel" about it.

OpenStudy (anonymous):

@Mertsj soo you think its my best bet to put down a fourth degree polynomial?

OpenStudy (mertsj):

I think you should ask someone in whom you have confidence.

OpenStudy (anonymous):

okay